A circle passes through the points A(-1,2),B (3,4) C(2,3).determine it's equation
1
Expert's answer
2020-08-17T19:31:07-0400
"Solution"

Start with the General Form of circle.

Substitution of A, B, C.


Let the circle be "x^2+y^2+Dx+Ey+F=0"


"(-1,2),\ (-1)^2+2^2-D+2E+F=0\\
(3,4),\ 3^2+4^2+3D+4E+F=0\\
(2,3),\ 2^2+3^2+2D+3E+F=0"


Simplification

"D-2E-F=5...........(1)\\
3D+4E+F=-25.....(2)\\
2D+3E+F=-13.....(3)"


Solve the equations.

"2(1)+(2),\ D=-3\\
(2)+(3),\ D+E=-12\\
(-3)+E=-12 \implies E=-12+3=-9\\
(-3)-2(-9)-F=5\\
\implies 15-F=5\\
\implies -F=-10 \implies F=10"


Substitute D, E, F in the General Form.


The equation of the circle is "x^2+y^2-3x-9y+10=0"